# Semiclassical treatment of bottomonium suppression and regeneration in $p+{\rm Pb}$ collisions

### v1

@@ III.4 Feed down contribution
As the bottomonia states propagate out of the hot QGP, the excited states can decay to lower-lying states, feeding these channels. To account for feed down, we define a matrix ($F$), which empirically relates the experimentally observed cross sections ($\vec{\sigma}_{\text{exp}}$) to the directly produced cross sections ($\vec{\sigma}_{\text{direct}}$) in $p+p$ collisions, $\vec{\sigma}_{\text{exp}}=F\vec{\sigma}_{\text{direct}}$. The quantities $\vec{\sigma}_{\text{direct}}$ and $\vec{\sigma}_{\text{exp}}$ contain the scattering cross sections for the bottomonia states $\Upsilon$(1S), $\Upsilon$(2S), $\chi_{b0}$(1P), $\chi_{b1}$(1P), $\chi_{b2}$(1P), $\Upsilon$(3S), $\chi_{b0}$(2P), $\chi_{b1}$(2P), and $\chi_{b2}$(2P) states before and after feed down, respectively. Further details on calculating $F$ are provided in Refs. [97, 65]. Incorporating feed down contributions into Eq. (16), the nuclear modification factor $R_{\rm pA}$ in min-bias $p+{\rm Pb}$ collisions for a given bottomonium state $i$ is expressed as:
@@
III.5 CNM and Primordial Suppression Results
We now present the combined results of primordial suppression with the total cold nuclear matter effects calculated in Sec. II. The total nuclear modification factor, including both the CNM and HNM effects, is
@@ III.5 CNM and Primordial Suppression Results
where the suppression due to CNM alone is given in Eq. (9). The factor $R^{\rm HNM}_{pA}$ is the primordial suppression factor due to the hot QGP medium calculated earlier in this section.
@@
The TAMU calculations employ the masses given in Table 2. Spin-weighted averages are used for the $\chi_{b}$ states: 9.8881 GeV and 10.2523 GeV for the $\chi_{b1}$(1P) and $\chi_{b2}$(2P) states respectively. Recall that there is no distinction between $\Upsilon$(3S) and $\chi_{b}$(2P) states for TAMU-P while the two are treated separately in the TAMU-NP calculations, compare Figs. 3 and 5.
@@ IV.3.1 $\Upsilon$(nnS) regeneration
Figures 13 and 14 illustrate our results as a function of $p_{T}$. Figure 13 gives the midrapidity results, $|y|<1.93$, for $\sqrt{s_{NN}}=5.02$ TeV compared to ATLAS [70] and CMS [71] measurements. The effect of regeneration is largest at low $p_{T}$ and decreases as $p_{T}$ increases. The regeneration contribution compensates for the substantial primordial suppression produced by large TAMU-NP reaction rates, filling the gap between the calculations with suppression alone and the data.
@@ V Conclusions and outlook
We have performed a systematic study of bottomonium suppression in minimum-bias $p+{\rm Pb}$ collisions at $\sqrt{s_{NN}}=5.02$ and 8.16 TeV. Our analysis incorporates both cold and hot nuclear matter effects. The cold matter effects include modifications of the nuclear parton distributions, coherent energy loss, and momentum broadening. The hot matter effects arise due to final-state interactions in the quark-gluon plasma. We have employed two complementary approaches to model the QGP-induced suppression: the KSU-QTraj approach, based on open quantum systems, and the semiclassical TAMU approach. The dissociation rates in the TAMU framework include predictions based on the $T$-matrix $U$ potential with perturbative medium coupling (TAMU-P) and fully integrated nonperturbative $T$-matrix calculations with self-consistent off-shell spectral functions and constrained by recent Wilson line correlators (TAMU-NP). We have also studied the effect of bottomonium regeneration in $p+{\rm Pb}$ collisions for the first time.
@@ Acknowledgements.
This work was supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through the Topical Collaboration in Nuclear Theory on Heavy-Flavor Theory (HEFTY) for QCD Matter under contract number DE-SC0023547. S.T. was supported by Kent State University through index 201452 and the Topical Collaboration in Nuclear Theory on Heavy-Flavor Theory (HEFTY) for QCD Matter under award no. DE-SC0023547. R.V. was also supported by the U.S. Department of Energy by LLNL under contract DE-AC52-07NA27344 and by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics (Nuclear Theory) under contract number DE-SC-0004014. R.R. and B.W. were also supported by the U.S. National Science Foundation under grant no. PHY-2209335. We would also like to thank Dr. Andrew Hanlon and Dr. Michael Strickland for valuable discussions on various aspects of this work.

### v2

@@ III.4 Feeddown contribution
As the bottomonium states propagate out of the hot QGP, the excited states can decay to lower-lying states, feeding these channels. To account for feeddown, we define a matrix ($F$), which empirically relates the experimentally observed cross sections ($\vec{\sigma}_{\text{exp}}$) to the directly produced cross sections ($\vec{\sigma}_{\text{direct}}$) in $p+p$ collisions, $\vec{\sigma}_{\text{exp}}=F\vec{\sigma}_{\text{direct}}$. The quantities $\vec{\sigma}_{\text{direct}}$ and $\vec{\sigma}_{\text{exp}}$ contain the production cross sections for the bottomonium states $\Upsilon$(1S), $\Upsilon$(2S), $\chi_{b0}$(1P), $\chi_{b1}$(1P), $\chi_{b2}$(1P), $\Upsilon$(3S), $\chi_{b0}$(2P), $\chi_{b1}$(2P), and $\chi_{b2}$(2P) states before and after feeddown, respectively. Further details on calculating $F$ are provided in Refs. [102, 65]. Incorporating feeddown contributions, the nuclear modification factor $R_{\rm pA}$ in min-bias $p+{\rm Pb}$ collisions for a given bottomonium state $i$ is expressed as:
@@
III.5 CNM and Primordial Suppression
We now present the combined results of primordial suppression with the total cold nuclear matter effects calculated in Sec. II. The total nuclear modification factor, including both the CNM and HNM effects on the inclusive $R^{\Upsilon}_{pA}$, defined in Eq. (15), is
@@ III.5 CNM and Primordial Suppression
where the suppression due to CNM alone is given in Eq. (9). The factor $R^{\rm HNM}_{pA}$ is the primordial suppression factor due to the hot QGP medium calculated earlier in this section. We note that because the CNM effects are independent of the final bottomonium state in our approach, feeddown only affects the primordial suppression in Eq. (16).
@@
The TAMU calculations employ the masses given in Table 2. Spin-weighted averages are used for the $\chi_{b}$ states: 9.8881 GeV and 10.2523 GeV for the $\chi_{b1}$(1P) and $\chi_{b2}$(2P) states respectively. Recall that there is no distinction between $\Upsilon$(3S) and $\chi_{b}$(2P) states for TAMU-P while the two are treated separately in the TAMU-NP calculations, compare Figs. 3 and 6.
@@ IV.3.1 $\Upsilon$(nnS) regeneration
Figures 14 and 15 illustrate our results as a function of $p_{T}$. Figure 14 gives the midrapidity results, $|y|<1.93$, for $\sqrt{s_{NN}}=5.02$ TeV compared to ATLAS [70] and CMS [71] measurements. The effect of regeneration is largest at low $p_{T}$ and decreases as $p_{T}$ increases. The regeneration contribution compensates for the substantial primordial suppression produced by large TAMU-NP reaction rates, filling the gap between the calculations with suppression alone and the data.
@@ V Conclusions and outlook
We have performed a systematic study of bottomonium modification in minimum-bias $p+{\rm Pb}$ collisions relative to $p+p$ collisions at $\sqrt{s_{NN}}=5.02$ and 8.16 TeV. Our analysis incorporates both cold and hot nuclear matter effects. The cold matter effects include modifications of the nuclear parton distributions, coherent energy loss, and momentum broadening. The hot matter effects arise due to final-state interactions in the quark-gluon plasma. We have employed two complementary approaches to model the QGP-induced suppression: the KSU-QTraj approach, based on open quantum systems, and the semiclassical TAMU approach. The dissociation rates in the TAMU framework include predictions based on the $T$-matrix $U$ potential with perturbative medium coupling (TAMU-P) and fully integrated nonperturbative $T$-matrix calculations with self-consistent off-shell spectral functions and constrained by recent Wilson line correlators (TAMU-NP). We have also studied the effect of bottomonium regeneration in $p+{\rm Pb}$ collisions for the first time.
@@ Acknowledgements.
This work was supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics through the Topical Collaboration in Nuclear Theory on Heavy-Flavor Theory (HEFTY) for QCD Matter under contract number DE-SC0023547. S.T. was supported by Kent State University through index 201452 and the Topical Collaboration in Nuclear Theory on Heavy-Flavor Theory (HEFTY) for QCD Matter under award no. DE-SC0023547. R.V. was also supported by the U.S. Department of Energy by LLNL under contract DE-AC52-07NA27344 and by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics (Nuclear Theory) under contract number DE-SC-0004014. R.R. and B.W. were also supported by the U.S. National Science Foundation under grant nos. PHY-2209335 and PHY-2514775. We would also like to thank Dr. Andrew Hanlon and Dr. Michael Strickland for valuable discussions on various aspects of this work.